OpenStudy (anonymous):

Evaluate 8 x minus eleven for x equals sixteen. nineteen forty eighty-eight one hundred seventeen

OpenStudy (phi):

not sure what you mean 8 x minus eleven for x equals sixteen. I get 8x - 11 =16 what does "for x" mean?

OpenStudy (anonymous):

tymes i guess

OpenStudy (accessdenied):

When I previously replied to this problem, I assumed that you were supposed to evaluate the expression, "8x-11" for x = 16.

OpenStudy (anonymous):

It means 8x - 11 if x = 16. Reinsert 16 for x: 8(16) - 11/128 - 11/117. 117 is the answer.

OpenStudy (anonymous):

ite

OpenStudy (anonymous):

117 would be the answer.